Honda purchases components from three suppliers. Components purchased from Supplier A are priced at $3 each and used at the rate of 18,900 units per month. Components purchased from Supplier B are priced at $5 each and used at the rate of 2,800 units per month. Components purchased from Supplier C are priced at $6 each and used at the rate of 1,300 units per month. The trucking company charges a fixed cost of $500 for the truck with an additional charge of $100 for each stop. Thus, if Honda asks for a pickup from only one supplier, the trucking company charges $500 + $100 = $600; from two suppliers, it charges $500 + 2*$100 = $700; from three suppliers, it charges $500 + 3*$100 = $800. Honda has a holding cost of 20%. Currently Honda purchases a separate truckload from each supplier. a) Calculate the optimal order quantity, cycle inventory, and total annual cost (excluding material cost) of each component at Honda with the current strategy. b) As part of its JIT drive, Honda has decided to aggregate purchases from the three suppliers. Calculate the optimal order quantity and cycle inventory of each component, and total annual cost (excluding material cost) for all three components at Honda with the complete aggregation strategy. (Please carry two decimal places for ?∗.) c) Which strategy is better, current strategy or complete aggregation strategy?
Answer #1

Note that 20% carrying cost is per month or per annum is not given. So, we have assumed it to be per annum. Eventually, we have converted all the monthly demand into annular demand by multiplying by 365.

(a)

Individual ordering strategy:

Common Fixed cost (S) = $500
Supplier (j) Annual Demand (Dj) Unit cost
(C
j)
Unit carrying cost
(H
j= 20% * Cj)
Individual ordering cost (Sj) EOQ
Q
j= (2.Dj.(S+Sj) / Hj)1/2
A 6,898,500 $3.0 $0.60 $100 117,461
B 1,022,000 $5.0 $1.00 $100 35,020
C 474,500 $6.0 $1.20 $100 21,784
Supplier (j) No of orders
N
j= Dj / Qj
Annual ordering cost = Nj.(S +Sj) Annual carrying cost = (Qj.Hj)/2 Total annual cost for Item-j
A 58.73 $35,238.08 $35,238.3 $70,476.38
B 29.18 $17,509.99 $17,510.0 $35,019.99
C 21.78 $13,069.23 $13,070.4 $26,139.63
Total cost $131,636.00

(b)

Complete aggregation:

Compute the common ordering frequency (n*) as follows:

n* = SQRT(i.(CA.DA+CB.DB.CC.DC)/(2*(S+SA+SB+SC)))) = SQRT(0.2*(3*6898500+5*1022000+6*474500)/(2*(500+100+100+100))) = 59.85

Common Fixed cost (S) = $500
Supplier (j) Annual Demand (Dj) Unit cost
(C
j)
Unit carrying cost
(H
j= 20% * Cj)
Individual ordering cost (Sj) n* Qj* = Dj / n*
A 6,898,500 $3.0 $0.60 $100 59.85 115263
B 1,022,000 $5.0 $1.00 $100 59.85 17076
C 474,500 $6.0 $1.20 $100 59.85 7928
Supplier (j) Annual ordering cost = n*.(S+∑jSj) Annual carrying cost = (Qj.Hj)/2 Total annual cost
A $47,880.00 $34,578.90
B - $8,538.00
C - $4,756.80
Totals $47,880.00 $47,873.70 $95,753.70

(c)

The second strategy is better for the lower total cost.